# Can Language Models Recognize Convincing Arguments?

### v1

@@ Abstract
The remarkable and ever-increasing capabilities of Large Language Models (LLMs) have raised concerns about their potential misuse for creating personalized, convincing misinformation and propaganda. To gain insights into LLMs’ persuasive capabilities without directly engaging in experimentation with humans, we propose studying their performance on the related task of detecting convincing arguments. We extend a dataset by Durmus & Cardie (2018) with debates, votes, and user traits and propose tasks measuring LLMs’ ability to (1) distinguish between strong and weak arguments, (2) predict stances based on beliefs and demographic characteristics, and (3) determine the appeal of an argument to an individual based on their traits. We show that LLMs perform on par with humans in these tasks and that combining predictions from different LLMs yields significant performance gains, even surpassing human performance. The data and code released with this paper contribute to the crucial ongoing effort of continuously evaluating and monitoring the rapidly evolving capabilities and potential impact of LLMs.
@@ 1 Introduction
As LLMs rise in capacity and popularity, so has the concern that they may enable the creation of tailor-made, convincing arguments capable of pushing false or misleading narratives online (De Angelis et al., 2023; Buchanan et al., 2021). While the dissemination of tailor-made misinformation predates the rise of LLMs (DiResta et al., 2019), the increasing accessibility and capability of state-of-the-art models such as GPT-4 and Claude could exacerbate this issue by enabling the creation of highly personalized misleading content with unprecedented ease (Bommasani et al., 2021; Goldstein et al., 2023).
Closely related to the task of generating persuasive content is the task of detecting content that would be particularly persuasive to individuals with specific demographics or beliefs. Malicious actors might not need to create content themselves; instead, they may promote existing content to specific users (Broniatowski et al., 2018). Furthermore, both tasks require that the underlying (large language) model encodes correlations between demographics and beliefs (e.g., U.S. Democrats tend to favor legalizing abortion), as well as the correlation between beliefs (e.g., anti-abortion individuals are more likely to be against gun control). Empirical work suggests that generation is easier than detection for LLMs (West et al., 2023), yet the latter task is better suited to continuously benchmark models’ persuasiveness capacities. Assessing models’ capacity to generate arguments requires continuous human experimentation as LLMs evolve, which can be time-consuming and resource-intensive. On the contrary, measuring a model’s capacity to detect content persuasive to specific demographics can be done quickly and without interaction with human subjects, making it a more efficient approach for benchmarking the persuasive capabilities of LLMs over time.
Figure 1: Our approach to study LLMs’ persuasiveness capabilities. We measure to which extent LLMs can reproduce human judgments on the quality and persuasiveness of arguments. If LLMs can predict users’ positions on stances (e.g., The death penalty should be legal) before and after reading a debate, as well as judge who was the better debater, we argue it would be well suited to generate personalized misinformation and propaganda.
@@
RQ1: judge the quality of arguments and identify convincing arguments as well as humans?
@@
Answering these questions can help assess the threat of LLM-generated misinformation and propaganda. If LLMs can detect good arguments (RQ1), determine the correlation between demographics and previously stated beliefs with people’s stances on new specific topics (RQ2), and determine whether an argument will convince specific individuals (RQ3), they are likely better at generating misinformation and propaganda.
To investigate these research questions, we extended a dataset collected by Durmus & Cardie (2018) from a defunct debate platform (debate.org). We annotate 833 politics-related debates with clear propositions, such as “The electoral college should remain unchanged.” Each debate contains arguments for (“Pro”) and against (“Con”) the proposition, along with votes from debate.org participants indicating the winning side. Importantly, the dataset includes demographic information of the voters as well as their stances on 48 so-called “big issues”. For 121 debates with 751 votes on three of the most prominent topics in the dataset, we obtained crowdsourced labels to compare the capabilities of LLMs to those of humans. Then, Using this enriched dataset, we evaluate the performance of four LLMs (GPT-3.5, GPT-4, Llama 2, and Mistral 7B) on three tasks: 1) identifying the side with more convincing arguments (RQ1); 2) predicting individuals’ stances on specific propositions before the debate, given their demographic and basic belief information (RQ2); and 3) predicting individuals’ stances on specific propositions after the debate, given their demographic and basic belief information (RQ3). Figure 1 illustrates our overall approach.
Our key finding is that LLM exhibits human-like performance across the three proposed tasks. In judging the better debater (RQ1), GPT-4 (Accuracy: 60.50%; RQ1) is as good as an individual voter in the dataset (Accuracy: 60.69%; RQ1). When predicting users’ stances on specific issues before and after reading the debate (RQ2 and RQ3), LLMs again perform similarly to humans. For instance, in the before the debate scenario (RQ2), Mistral yields an accuracy of 42.27%, whereas crowdworkers achieve 39.86% (random guessing would yield 33.3% accuracy). However, zero-shot prediction with LLMs still underperforms a supervised machine learning model [XGBoost (Chen & Guestrin, 2016)], which achieves 58.25% accuracy in cross-validation. Nevertheless, stacking the predictions of LLMs and using them as features in a supervised learning setting reduces the performance gap (45.91%).
Overall, our work contributes to the growing body of research on the societal impact of LLMs (Bommasani et al., 2021; Solaiman et al., 2023; Weidinger et al., 2023, inter alia). By investigating LLMs’ ability to detect persuasive content tailored to specific demographics, we shed light on their potential for generating targeted misinformation and propaganda.
@@ 2 Related Work
We review related work in three broad directions closely related to the task.
@@
Demographics have long been known to impact people’s political beliefs and attitudes. Group-level demographic factors such as race, religion, and education shape individuals’ perspectives on various political issues and voting behavior in the U.S. context (Campbell et al., 1960; Erikson & Tedin, 2019). For example, 78% of Black, 72% of Asian, and 65% of Hispanic workers see efforts on increasing diversity, equity, and inclusion at work positively, compared to 47% of White workers (Minkin, 2023). Similarly, previous work indicates that persuasion depends on the message recipients’ existing values and that individual differences can influence persuasion (O’Keefe, 2015). For instance, Hirsh et al. (2012) showed that tailoring messages to different personality traits can make them more persuasive; Orji et al. (2015) showed that men and women differ significantly in their responsiveness to the different persuasive strategies. And, closer to the work at hand, Durmus & Cardie (2018) and Al Khatib et al. (2020) have shown that modeling or considering debate characteristics can enhance the prediction of argument persuasiveness. In this work, we examine how LLMs can capture the correlations between demographics and beliefs (RQ2) and how personal traits determine the persuasiveness of arguments (RQ3).
Argument quality.
Defining argument quality is no easy task, or as persuasion scholars O’Keefe & Jackson (1995) put it: “there is no clear general abstract characterization of what constitutes argument quality.” For example, an argument may be good because of its effectiveness, i.e., in practice, it convinces people (O’Keefe & Jackson, 1995), or because its congency, i.e., that the argument is composed of individually accepted premises that allow its conclusion to be reached (Johnson & Blair, 2006), or even because it is reasonable, i.e., the argument and the way it is put is acceptable contributes to resolving the disagreement (Walton, 2005). Over the last decades, there’s been excitement in the NLP and artificial intelligence communities about automatically extracting arguments from text (Lawrence & Reed, 2020; Cabrio & Villata, 2018) (or even multimodal data, see Liu et al. (2023)) and using computational methods to better understand what makes arguments effective (Wachsmuth et al., 2017; Tan et al., 2016). In this work, we examine to which extent LLMs can capture which debates were considered to have better arguments (RQ1) and to which extent arguments were effective across different demographics (RQ3).
@@
Microtargeting or “personalized persuasion” refers to tailoring the language or content of messages to individuals based on their characteristics (e.g., demographics and prior political beliefs) to make them maximally persuasive. Evidence on the effect sizes of microtargeting is mixed (Guess & Coppock, 2020; Coppock et al., 2020; Matz et al., 2017; Tappin et al., 2023), which has led Teeny et al. (2021) to propose that research on microtargeting should move from “does microtargeting work?”, to “when micro-targeting works?” At the same time, the increasing popularity and capabilities of large language models (LLMs) have raised concerns that they may make microtargeting cheaper and better and that they may enable new ways to “microtarget” misinformation and propaganda, e.g., personalized chatbots – see  Goldstein et al. (2023) for a comprehensive discussion. This concern has been corroborated by new work suggesting that LLMs are capable of generating messages perceived as equally or more persuasive than humans (Bai et al., 2023) that they can personalize messages to make them more persuasive (Simchon et al., 2024) and that LLMs can successfully persuade humans in debates by exploiting their personal traits Salvi et al. (2024). This work is primarily motivated by the theme of microtargeting, as our goal is to continuously benchmark LLMs capabilities of tailor messages so that they are persuasive to different individuals (RQ3).
3 Data
Data for this study was collected by Durmus & Cardie (2018) from an online debate platform (debate.org; no longer operational). The platform allowed users to participate in and vote on debates covering a breadth of topics, including politics, religion, and science. Each debate within the dataset consists of multiple rounds, each round with an argument from both the “Pro” and “Con” perspectives. Users on the platform could vote on various aspects of the debate, such as which side they believed provided a more convincing argument. The raw dataset contains 78,376 debates, 45,348 users, and 195,724 votes. Each user has corresponding demographic information, such as gender and age, as well as their stances on 48 so-called “big issues,” such as abortion, capital punishment, and national health care. Nevertheless, most demographic data is missing from the dataset. Most important to the work at hand, voters had to indicate which side: 1) Made more convincing arguments; 2) They agreed with before the debate; 3) They agreed with after the debate. We measure LLMs’ capacity to recognize convincing arguments by predicting the responses to these three questions, each of which could be answered “Pro,” “Con,” or “Tie.” Note that predicting question #1–#3 corresponds to our research questions RQ1RQ3.
@@ 3 Data
PoliProp [PP]. We study these 833 annotated debates, considering all votes ($n=4$,$871$) in these debates for users with no more than five missing values in demographic information ($4$,$871$ out of $7{,}797$). We also trimmed each debate in the dataset larger than the smallest context window (4096 tokens) among LLMs considered. Trimming is done by removing one round at a time from the end of the debate until the token count is small enough, an approach that equally penalizes both debaters (unlike simply removing tokens at the end of the debate). Hereafter, we call this dataset the PoliProp dataset.
PoliIssues [IS]. We also separately consider all debates on abortion ($n=50$), gay marriage ($n=51$), and capital punishment ($n=31$), the most prominent topics in the dataset. Given that debates within the three themes are similar, we use this data to compare LLM performances with traditional machine learning methods, predicting participants’ votes using their demographic and stances on big issues as features. To obtain a human baseline, we collect crowdsourced labels using Amazon Mechanical Turk (MTurk) for each of the 751 votes cast on these 121 debates. Crowdworkers are essentially presented with the same questions as the LLMs. Given a debate, we ask who gave the better arguments. Given a set of characteristics by a voter as well as the debate, we ask whether the voter would have agreed with the proposition before and after reading the debate. Hereafter, we call this dataset the PoliIssues dataset. We provide additional information on the crowdsourcing methodology in Appendix B.
@@ 4 Methods
For this study, we compare the performance of two open-source LLMs, namely Mistral 7B (Mistral-7b-Instruct-v0.1) and Meta’s Llama 2 70B (Llama-2-70b-chat), with OpenAI’s closed-source GPT-3.5 (gpt-3.5-turbo-1106) and GPT-4 (gpt-4-0613).
@@
We follow Staab et al. (2023) to develop our prompt: each had a system role, context, question, and constraint. We experimented with different structures and found that, overall, the structure mattered little as long as the wording was clear and concise. Since we had three research questions to answer, we had three prompt structures that combined the debate proposition, the debate itself, and user demographics. We show the prompt structure used for RQ1 in Figure 2. All prompts indicated that LLMs should respond only with the labels “Pro,” “Con,” or “Tie”. Nevertheless, many of the models failed to adhere to this instruction, necessitating post-processing to extract the actual answer. Generally, instances of incorrect responses involved the answer accompanied by additional spaces or punctuation or presented in a complete sentence format, such as: “Based on the given demographics, the person is most likely to agree with the ‘Con’ side in the debate.” We used heuristics to extract the responses in these cases. Additionally, there were occasions when the LLMs failed to produce any answer, resulting in responses akin to: “I cannot determine the person’s position in the debate without additional information.” We depict the remaining prompts and provide details in Appendix A.
@@
We evaluate the accuracy of language models by comparing the answers they provide with the ground truth data from PoliProp and PoliIssues. We obtain confidence intervals through bootstrapping. Besides considering each LLM individually, we also consider the performance of a “stacked” LLM, obtained by using the output of several large language models (with slightly different instantiations) as features in a supervised machine learning model.
@@
  • MTurk; (RQ2–RQ3) We crowdsourced the tasks for each research question for the PoliIssues dataset, obtaining a human equivalent answer to the questions we asked the LLMs. These are detailed in Appendix B.

### v2

@@ Abstract
The capabilities of large language models (LLMs) have raised concerns about their potential to create and propagate convincing narratives. Here, we study their performance in detecting convincing arguments to gain insights into LLMs’ persuasive capabilities without directly engaging in experimentation with humans. We extend a dataset by Durmus and Cardie (2018) with debates, votes, and user traits and propose tasks measuring LLMs’ ability to (1) distinguish between strong and weak arguments, (2) predict stances based on beliefs and demographic characteristics, and (3) determine the appeal of an argument to an individual based on their traits. We show that LLMs perform on par with humans in these tasks and that combining predictions from different LLMs yields significant performance gains, surpassing human performance. The data and code released with this paper contribute to the crucial effort of continuously evaluating and monitoring LLMs’ capabilities and potential impact. (https://go.epfl.ch/persuasion-llm)
Can Language Models Recognize Convincing Arguments?
Paula Dolores Rescala1, Manoel Horta Ribeiro1, Tiancheng Hu2, Robert West1 1 EPFL, 2 University of Cambridge Correspondence: manoel.hortaribeiro@epfl.ch
@@ 1 Introduction
As LLMs rise in capacity and popularity, so has the concern that they may help create and propagate tailor-made, convincing narratives De Angelis et al. (2023); Buchanan et al. (2021). While “tailor-made misinformation” predates LLMs (DiResta et al., 2019), frontier models such as GPT-4, Claude 3, and Gemini 1.5 could add to the problem by allowing malicious actors to easily create diverse, personalized content (Bommasani et al., 2021; Goldstein et al., 2023) or enable the detection (and amplification) of existing content that would be particularly persuasive to individuals with specific demographics or beliefs (Broniatowski et al., 2018).
Previous work has found LLMs to be persuasive in the generative setting Simchon et al. (2024); Hackenburg and Margetts (2024); Breum et al. (2024); for example, Salvi et al. (2024) found that, when provided with personal attributes, GPT-4 outperformed crowdworkers in a debate setting. Yet, assessing models’ capacity to generate arguments requires continuous human experimentation as LLMs evolve, which can be time-consuming and resource-intensive. On the contrary, measuring a model’s capacity to detect content persuasive to specific demographics can be done quickly and without interaction with human subjects, making it a more efficient approach for benchmarking the persuasive capabilities of LLMs.
Figure 1: Our approach to study LLMs’ persuasiveness capabilities. We measure to which extent LLMs can reproduce human judgments on the quality and persuasiveness of arguments. Suppose LLMs can predict users’ positions on stances (e.g., The death penalty should be legal) before and after reading a debate and judge who the better debater was. In that case, they would be well suited to power personalized misinformation and propaganda.
Present Work.
@@
  • RQ1: judge the quality of arguments and identify convincing arguments and humans?
@@
To investigate these questions, we extend a dataset collected by Durmus and Cardie (2018) from a defunct debate platform (debate.org). We annotate 833 politics-related debates with clear propositions, such as “The electoral college should remain unchanged.” Each debate contains arguments for (“Pro”) and against (“Con”) the proposition, along with votes from debate.org participants indicating the winning side. Importantly, the dataset includes demographic information of the voters as well as their stances on 48 so-called “big issues”. For 121 debates with 751 votes on three of the most prominent topics in the dataset, we obtained crowdsourced labels to compare the capabilities of LLMs to those of humans. Then, using this enriched dataset, we evaluate the performance of four LLMs (GPT-3.5, GPT-4, Llama 2, and Mistral 7B) on three tasks: 1) identifying the side with more convincing arguments (RQ1); 2) predicting individuals’ stances on specific propositions before the debate, given their demographic and basic belief information (RQ2); and 3) predicting individuals’ stances on specific propositions after the debate, given their demographic and basic belief information (RQ3). Figure 1 illustrates our approach.
Our key finding is that LLM exhibits human-like performance across the three proposed tasks. In judging the better debater (RQ1), GPT-4 (Accuracy: 60.50%) is as good as an individual voter in the dataset (Accuracy: 60.69%;). When predicting users’ stances on specific issues before and after reading the debate (RQ2 and RQ3), LLMs again perform similarly to humans. For instance, in the before the debate scenario (RQ2), Mistral yields an accuracy of 42.27%, whereas crowdworkers achieve 39.86% (random guessing would yield 33.3% accuracy). However, zero-shot prediction with LLMs still underperforms a supervised machine learning model [XGBoost (Chen and Guestrin, 2016)], which achieves 58.25% accuracy in cross-validation. Nevertheless, stacking the predictions of LLMs and using them as features in a supervised learning setting reduces the performance gap (45.91%).
Overall, our work contributes to the growing body of research on the societal impact of LLMs (Bommasani et al., 2021; Solaiman et al., 2023; Weidinger et al., 2023, inter alia). We shed light on the potential misuse of LLMs by investigating their ability to detect persuasive content tailored to specific demographics.
@@ 2 Related Work
We review related work in three broad directions closely related to the tasks proposed.
@@
Demographics have long been known to impact people’s political beliefs and attitudes. Group-level demographic factors such as race, religion, and education shape individuals’ perspectives on various political issues and voting behavior in the U.S. context (Campbell et al., 1960; Erikson and Tedin, 2019). For example, 78% of Black, 72% of Asian, and 65% of Hispanic workers see efforts on increasing diversity, equity, and inclusion at work positively, compared to 47% of White workers (Minkin, 2023). Similarly, previous work indicates that persuasion depends on the message recipients’ existing values and that individual differences can influence persuasion (O’Keefe, 2015). For instance, Hirsh et al. (2012) demonstrated that tailoring messages to different personality traits can make them more persuasive; Orji et al. (2015) showed that men and women differ significantly in their responsiveness to the different persuasive strategies. Closer to the work at hand, Durmus and Cardie (2018) and Al Khatib et al. (2020) showed considering demographic characteristics can enhance the prediction of argument persuasiveness. However, the extent to which LLMs can utilize demographic characteristics in persuasiveness judgment remains underexplored. In this work, we examine how LLMs can capture the correlations between demographics and beliefs (RQ2) and how personal attributes determine the persuasiveness of arguments (RQ3).
Argument Mining and Argument quality.
Defining argument quality is no easy task, or as persuasion scholars O’Keefe and Jackson (1995) have put it: “there is no clear general abstract characterization of what constitutes argument quality.” An argument may be deemed good due to its effectiveness in convincing people (O’Keefe and Jackson, 1995), its cogency from individually accepted premises that lead to a conclusion (Johnson and Blair, 2006), or its reasonableness in contributing to resolving a disagreement (Walton, 2005). Over recent decades, there has been significant interest in automatically extracting arguments from text (Habernal and Gurevych, 2016, 2017; Swanson et al., 2015, inter alia), as detailed in surveys like (Cabrio and Villata, 2018; Lawrence and Reed, 2020). Additionally, research has explored computational argument quality and persuasiveness analysis (Habernal and Gurevych, 2016; Tan et al., 2016; Wachsmuth et al., 2017, inter alia). Contemporary works begin to explore the potential of LLMs in argument quality judgement Mirzakhmedova et al. (2024); Wachsmuth et al. (2024). Our work complements existing work by examining the extent to which LLMs can identify higher-quality arguments in a debate setting (RQ1) and determine argument effectiveness across individuals with different demographics and beliefs (RQ3).
@@
Microtargeting or “personalized persuasion” refers to tailoring the language or content of messages to individuals based on their characteristics (e.g., demographics and prior political beliefs) to make them maximally persuasive. Evidence on the effect of microtargeting is mixed (Guess and Coppock, 2020; Coppock et al., 2020; Matz et al., 2017; Tappin et al., 2023), which has led Teeny et al. (2021) to propose that research on microtargeting should move from “does microtargeting work?”, to “when micro-targeting works?” At the same time, the increasing popularity and capabilities of LLMs have raised concerns that they may not only make microtargeting cheaper and more effective but also enable new forms of “microtargeting” misinformation and propaganda, such as through personalized chatbots—see Goldstein et al. (2023) for a comprehensive discussion. These concerns are corroborated by recent studies suggesting that LLMs are capable of generating messages perceived as equally or more persuasive than humans (Bai et al., 2023; Durmus et al., 2024); that they can personalize messages to make them more persuasive (Simchon et al., 2024); and that LLMs can successfully persuade humans in debates by exploiting their personal traits Salvi et al. (2024). One key drawback, though, is that these studies typically involve large and expensive experiments that cannot easily be replicated when a new LLM is released or explore the large hyperparameter space of existing models (e.g., prompting strategy and decoding algorithm). In our work, we argue that we could instead evaluate the effectiveness of LLM in determining whether someone of a specific set of demographic characteristics would find an argument convincing (RQ3) and view this as a proxy of the LLM’s ability to perform political microtargeting.
Social sensing.
Prediction tasks where individuals are asked to determine the preferences and opinions of others have been broadly referred to as social sensing Galesic et al. (2021). Previous work using this approach has shown that human social sensing outperformed traditional polling in forecasting elections Galesic et al. (2018) and that the approach is useful in predicting disease outbreaks Christakis and Fowler (2010). Here, the tasks associated with RQ2 and RQ3 are, in their essence, social sensing tasks, as we ask LLMs (and crowdworkers) about the preferences and opinions of others. Although informative, predictions obtained through human social sensing are known to be subjected to biases Ross et al. (1977); Chambers and Windschitl (2004), and therefore, it is possible that so are predictions obtained through LLM social sensing.
@@
US Hegemony U.S. hegemony is desirable.
Abortion Abortion should be illegal.
@@
3 Data
Data for this study was collected by Durmus and Cardie (2018) from an online debate platform (debate.org; no longer operational). The platform allowed users to participate in and vote on debates covering a breadth of topics, including politics, religion, and science. Each debate within the dataset consists of multiple rounds, each round with an argument from both the “Pro” and “Con” perspectives. Users on the platform could vote on various aspects of the debate, such as which side they believed provided a more convincing argument. The raw dataset contains 78,376 debates, 45,348 users, and 195,724 votes. Each user has corresponding demographic information, such as gender and age, as well as their stances on 48 so-called “big issues,” such as abortion, capital punishment, and national health care (see Appendix B for details). Nevertheless, most demographic data is missing from the dataset. Most important to the work at hand, voters had to indicate which side: 1) Made more convincing arguments; 2) They agreed with before the debate; 3) They agreed with after the debate. We measure LLMs’ capacity to recognize convincing arguments by predicting the responses to these three questions, each of which could be answered “Pro,” “Con,” or “Tie.” Note that predicting question #1–#3 corresponds to our research questions RQ1RQ3.
PoliProp [PP].
We study these 833 annotated debates, considering all votes ($n=4$,$871$) in these debates for users with no more than five missing values in demographic information ($4$,$871$ out of $7{,}797$). We also trimmed each debate in the dataset larger than the smallest context window (4096 tokens) among LLMs considered. Trimming is done by removing one round at a time from the end of the debate until the token count is small enough, an approach that equally penalizes both debaters (unlike simply removing tokens at the end of the debate). Hereafter, we call this the PoliProp dataset.
PoliIssues [IS].
We also separately consider all debates on abortion ($n=50$), gay marriage ($n=51$), and capital punishment ($n=31$), the most prominent topics in the dataset. Given that debates within the three themes are similar, we use this data to compare LLM performances with traditional machine learning methods, predicting participants’ votes using their demographic and stances on big issues as features. To obtain a human baseline, we collect crowdsourced labels using Amazon Mechanical Turk (MTurk) for each of the 751 votes cast on these 121 debates. Crowdworkers are essentially presented with the same questions as the LLMs. Given a debate, we ask who gave the better arguments. Given a set of characteristics by a voter as well as the debate, we ask whether the voter would have agreed with the proposition before and after reading the debate. Hereafter, we call this dataset the PoliIssues dataset. For more information on crowdsourcing, see Appendix C.
Figure 2: Prompt structure used in RQ1.
Figure 3: Prompt structure used in RQ2.
Figure 4: Prompt structure used in RQ3.
@@ 4 Methods
For this study, we compare the performance of two open-source LLMs, namely Mistral 7B (Mistral-7b-Instruct-v0.1) and Meta’s Llama 2 70B (Llama-2-70b-chat), with OpenAI’s closed-source GPT-3.5 (gpt-3.5-turbo-1106) and GPT-4 (gpt-4-0613). We use the standard temperatures for each model.
@@
We follow Staab et al. (2023) to develop our prompt: each had a system role, context, question, and constraint. We experimented with different structures and found that, overall, the structure mattered little as long as the wording was clear and concise. Since we had three research questions to answer, we had three prompt structures that combined the debate proposition, the debate itself, and user demographics. We show the prompt structure used for RQ1 in Figure 2. All prompts indicated that LLMs should respond only with the labels “Pro,” “Con,” or “Tie”. Nevertheless, many of the models failed to adhere to this instruction, necessitating post-processing to extract the actual answer. Generally, instances of incorrect responses involved the answer accompanied by additional spaces or punctuation or presented in a complete sentence format, such as: “Based on the given demographics, the person is most likely to agree with the ‘Con’ side in the debate.” We used heuristics to extract the responses in these cases. Additionally, there were occasions when the LLMs failed to produce any answer, resulting in responses akin to: “I cannot determine the person’s position in the debate without additional information.” We depict the remaining prompts in Figures 3 and 4 and provide further details in Appendix A.
@@
We evaluate the accuracy of language models by comparing the answers they provide with the ground truth data from PoliProp and PoliIssues. We obtain confidence intervals through bootstrapping. Besides considering each LLM individually, we also consider the performance of stacked LLM predictions, obtained by using the output of different LLMs as features in a supervised machine learning model Hastie et al. (2009).
@@
  • MTurk; (RQ2–RQ3) We crowdsourced the tasks for each research question for the PoliIssues dataset, obtaining a human equivalent answer to the questions we asked the LLMs. These are detailed in Appendix C.